At the beginning of each year, the Internal Revenue Service (IRS) releases information on the likelihood...

  1. At the beginning of each year, the Internal Revenue Service (IRS) releases information on the likelihood of a tax return being audited. In 2015, the IRS audited 1,228,117 individual tax returns from the total of 146,861,217 filed returns; also, the IRS audited 24,761 returns from the total of 1,915,337 corporation returns filed (IRS 2015 Data Book). Round all answers to four decimal places.
    1. Suppose an individual tax return from 2015 is randomly selected. What is the probability that the return was audited by the IRS?
    2. Refer to part a. Determine the probability that an individual return was not audited by the IRS.
    3. Suppose a corporation tax return from 2015 is randomly selected. What is the probability that the return was audited by the IRS?
    4. Refer to part c. Determine the probability that a corporation return was not audited by the IRS.
Answer #1
individual corporation total
audited 1228117 24761 1252878
not audited 145633100 1890576 147523676
total 146861217 1915337 148776554

a)

probability that the return was audited by the IRS =1228117/146861217 =0.0084

b) e probability that an individual return was not audited by the IRS =1-0.0084 =0.9916

c) probability that the return was audited by the IRS =24761/1915337 =0.0129

d) probability that a corporation return was not audited by the IRS =1-0.0129 =0.9871